Brandtly Industries invests a large sum of money in R&D; as a result, it retains and reinvests all of its earnings. In other words, Brandtly does not pay any dividends, and it has no plans to pay dividends in the near future. A major pension fund is interested in purchasing Brandtly's stock. The pension fund manager has estimated Brandtly's free cash flows for the next 4 years as follows: $4 million, $7 million, $11 million, and $16 million. After the fourth year, free cash flow is projected to grow at a constant 7%. Brandtly's WACC is 16%, the market value of its debt and preferred stock totals $52 million, the firm has $16 million in non-operating assets, and it has 25 million shares of common stock outstanding.

a. What is the present value of the free cash flows projected during the next 4 years? Do not round intermediate calculations. Round your answer to the nearest dollar. Write out your answers completely. For example, 13 million should be entered as 13,000,000.

24534308.15

b. What is the firm's horizon, or continuing, value? Round your answer to the nearest dollar. Write out your answers completely. For example, 13 million should be entered as 13,000,000.

190222222

c. What is the market value of the company's operations? Do not round intermediate calculations. Round your answer to the nearest dollar. Write out your answers completely. For example, 13 million should be entered as 13,000,000.

?

What is the firm's total market value today? Do not round intermediate calculations. Round your answer to the nearest dollar. Write out your answers completely. For example, 13 million should be entered as 13,000,000.

?

d. What is an estimate of Brandtly's price per share? Do not round intermediate calculations. Round your answer to the nearest cent.

?
Answer #1
Present Value(PV) of Cash Flow:
(Cash Flow)/((1+i)^N)
i=discount rate =WACC=`16%=0.16
N=Year   of Cash Flow
a Present Value of Free Cash Flows
N Year 1 2 3 4
FCF Free Cash Flows $4,000,000 $7,000,000 $11,000,000 $17,000,000 SUM
PV=FCF/(1.16^N) Present Value of Free Cash Flows $3,448,275.86 $5,202,140.31 $7,047,234.41 $9,388,948.66 $25,086,599.24
Present Value of Free Cash Flows during next4 years       25,086,599.24
b Horizon or continuing Value
FCF in year 4 $17,000,000
g=Growth rate in year 5 onwards=7% 0.07
FCF in year 5=17 million*(1+g)= $18,190,000 (17000000*1.07)
Required rate of return =R=16% 0.16
Horizon or continuing Value $202,111,111.11 (18190000/(R-g)= (18190000/(0.16-0.07)
Horizon or continuing Value     202,111,111.11
c Market Value of company's operations
A Present Value of Horizon $111,624,167.45 (202111111.11/(1.16^4)
B Present Value of Year1 to 4 FCF       25,086,599.24
C=A+B Total Enterprise Value $136,710,766.69
D Less: Non Operating Assets $16,000,000
E=C-D Market Value of company's operations $120,710,766.69
Total Market Value today $136,710,766.69
d Price per share
F Total Market Value today $136,710,766.69
G Market Value of Debt and preferred stock $52,000,000
H=F-G Market Value of Common Equity $84,710,766.69
I Number of shares of common stock outstanding             25,000,000
J=H/I Estimated Price per share $3.3884307
Estimated Price per share $3.39